### v2

Abstract
The rapid advancement of generative artificial intelligence has positioned text-to-video (T2V) models as essential components for building future multimodal world simulators. However, current T2V systems remain highly susceptible to jailbreak attacks, where carefully crafted prompts circumvent safety mechanisms and induce the generation of harmful or unsafe content. Such vulnerabilities severely undermine the reliability and security of simulation-driven applications. In this work, we present T2VShield, a comprehensive and model-agnostic defense framework designed to safeguard T2V models against jailbreak attacks. Our approach first systematically examines the three critical stages of T2V security (i.e., input, model, and output stages) and reveals inherent limitations in existing defenses, including (i) prompt structures that exploit semantic ambiguities, (ii) the challenge of detecting malicious content in high-dimensional and temporally dynamic outputs, and (iii) the rigidity of model-centric mitigation strategies. Based on the observations, T2VShield integrates an input rewriting mechanism based on Chain-of-Thought (CoT) reasoning and multimodal GraphRAG retrieval, enabling semantic sanitization of malicious prompts at the input level. Additionally, we introduce a multi-scope output detection module to capture both local anomalies and global inconsistencies from multi-timescale slicing and multimodal feature fusion, ensuring robust, time-aware defense. Crucially, T2VShield is plug-and-play, requiring no access to internal model parameters and supporting both open- and closed-source T2V systems. Extensive evaluations across two open-source and three commercial T2V platforms show that T2VShield reduces jailbreak success rates by up to 35% compared to state-of-the-art baselines. Furthermore, we design a human-centered audiovisual evaluation protocol to assess perceptual safety, highlighting the critical role of visual defense in enhancing the trustworthiness of next-generation multimodal simulators.
keywords:
Text-to-Video Model, Safety Evaluation Framework, Jailbreak Defense
[…]
1 Introduction
The rapid advancement of generative artificial intelligencekong2024hunyuanvideo ; blattmann2023stable ; kondratyuk2023videopoet ; singer2022make ; wu2023tune ; zhang2024show is steering the field toward modeling complex, multimodal worlds, with visual and auditory perception playing central roles mclachlan2010central ; qin2024worldsimbench . Among these, text-to-video (T2V) models have emerged as a critical component in the development of world simulators, leveraging the primacy of visual information in human cognition (estimated to account for over 80% of external perception) and the technological maturity of unimodal generative methods, such as diffusion models and Transformer-based architectures. Recent state-of-the-art T2V systems (e.g., Sora liu2024sora , OpenSora zheng2024open , and CogVideoX yang2024cogvideox ) aim to synthesize video sequences from natural language descriptions with high temporal coherence, visual realism, and physical plausibility. This task demands not only the generation of photorealistic individual frames but also continuity in motion, consistency in object appearance and positioning, and accurate modeling of real-world physical dynamicssun2024sora ; nair2023t2v ; li2024t2v ; tian2024videotetris ; liao2024evaluation .
However, recent research efforts miao2024t2vsafetybench ; yoonsafree have revealed that both open-source and closed-source T2V systems are susceptible to jailbreak attacksxie2023defending ; li2024semantic ; yu2024don ; jin2024attackeval , whereby adversaries ying2024jailbreak ; chu2024comprehensive ; madry2017towards ; akhtar2021advances craft sophisticated prompts to bypass safety mechanisms and induce harmful or non-compliant video generation. As shown in Fig. 1, these vulnerabilities compromise the safety of the models themselves and threaten the broader trustworthiness of multimodal world simulators, particularly in applications involving media generation, education, and virtual environments. In this study, we propose T2VShield, a comprehensive and model-agnostic defense framework designed to safeguard T2V models against jailbreak attacks.
Our framework systematically examines defense mechanisms across all stages of deployment (i.e., input, model, and output stages), offering both a conceptual and practical blueprint for robust protection. Through empirical analysis, we uncover key challenges that undermine existing defenses. (1) Prompt-based vulnerabilities: Jailbreak prompts often reveal semantic cues or structural patterns that can be parsed or filtered, but current systems underexploit this signal. (2) Temporal sparsity and randomness: Harmful content may emerge transiently within a single frame or across disjoint moments in a video, making detection based on frame-by-frame analysis unreliable. (3) Model-centric constraints: Although training-based defense to T2V models may improve robustness in theory, they are typically costly, complex, and disruptive to the generation quality of benign inputs, making them unsuitable for practical deployment.
Building on these, we introduce a dual-phase framework that prevent malicious input. At the input level, we combine CoT-based reasoning with multimodal example retrieval (PosNegRAG) to improve the model’s ability to identify and rewrite risky prompts. At the output level, we propose a multiscale detection framework that integrates frame-level vision cues with global semantic summaries to capture both local and global risks. Our defense mechanism is model-agnostic, requiring no access to internal parameters, and can be seamlessly integrated with different T2V systems.
Figure 1: Jailbreak attack and defense against text-to-video models. The top figure illustrates an attack scenario where various jailbreak hints lead to the generation of insecure videos for both open-source and closed-source T2V models. The bottom figure shows the defense mechanisms, including input-level, model-level, and output-level defenses, used to mitigate the effects of jailbreak attacks and generate sanitized videos.
To demonstrate the efficacy, we conduct extensive experiments over 2 benchmarks on 2 open-source and 3 closed-source T2V models, reduces jailbreak success rates by up to 35% compared to state-of-the-art baselines. In addition, we design complementary validation experiments based on human perception to jointly assess the visual and auditory perceptual security of generated videos, further validating the effectiveness of our defense mechanism in mitigating potentially harmful outputs. Experimental results highlight that securing video frame content is fundamental to enhancing the overall trustworthiness of future multimodal world simulator systems. Our main contributions can be summarized as follows:
  • We propose the first systematic jailbreak defense framework for T2V models, which encompasses multi-stage security evaluation across the input, model, and output stages. Based on thorough examination, we propose inherent limitations in existing defenses, including the structural vulnerabilities of attack prompts, the challenge posed by high-dimensional video dynamics for output detection, and the practical limitations of model-side defenses.
  • We propose T2VShield innovatively introduces an input rewriting strategy based on CoT inference guided by PosNegRAG retrieval and a multi-scope output detection module based on multi-timescale slicing and multimodal feature fusion , improving accuracy and enabling seamless deployment across both open-source and closed-source T2V systems to significantly enhance content safety.
  • We comprehensively evaluate our defense on two open-source models (OpenSora zheng2024open , CogVideoX yang2024cogvideox ) and three closed-source models (Sora liu2024sora , Kling kling2024 , Pika pika2024 ) where we can reduce the attack success rate of the commercial model by an average of 33% to a level where the ASR does not exceed 15%., and further validate the necessity of visual-level defenses through audiovisual experiments.
2 Related Work
2.1 Video Generation with Text Condition
T2V is a technique for generating dynamic video content directly from natural language prompts. In recent years, T2V technology has made rapid progress with the development of generative models. Early approaches such as Video Diffusion Models (VDM) ho2022video , inspired by traditional image diffusion network (e.g., U-Net ronneberger2015u ) architectures, explored the initial paths of text-to-video generation by training on joint images and short video clips.
With the advent of the diffusion model era, a series of models based on DiT architectures began to receive widespread attention. Representative works include: Open-Sora zheng2024open is one of the earliest community-released T2V models based on a DiT-style backbone, enabling high-quality, long-duration video generation with large-scale computational support. CogVideoX yang2024cogvideox , on the other hand, introduces a 3D convolution mechanism to better model spatiotemporal features, enabling the generation of long videos with strong temporal coherence, while VideoCrafter chen2023videocrafter1 is optimized based on the Stable Diffusion architecture through fine-tuning, and generates high-quality short videos with artistic style adaptability under lightweight deployment conditions.
Meanwhile, along with the popularity of short videos and the growing demand for commercialization, closed-source T2V systems have significantly improved in terms of generation quality. The Sora model demonstrates for the first time the ability to generate videos that follow the rules of the physical world, with long durations and high fidelity. Pika proposes a lightweight video generation tool for creative workers, supporting a variety of output styles such as 3D animation, cinematography, and hand-drawn styles. Luma luma2024 combines NeRF-based 3D reconstruction with multimodal video generation technologies, simulating realistic light and shadow interaction effects, while Kling kling2024 supports narrative video generation with multi-character interactions, especially excelling in Chinese film, television, and ancient-style storytelling.
However, studies such as T2VSafetyBench miao2024t2vsafetybench reveal that both open-source and closed-source T2V systems are generally at risk of jailbreak attacks, with a higher likelihood of generating potentially harmful content. Such vulnerabilities not only seriously impact the reputation of commercial services but also threaten the credibility of world simulators.
2.2 Jailbreak Attack and Defense
Intelligent systems are vulnerable to attacks, such as adversarial examples liu2019perceptual ; liu2020bias ; liu2020spatiotemporal ; liu2023x ; liu2023exploring ; liu2021training ; wei2018transferable , data poisoning LiuCompromising2024 ; liu2025pre ; liang2024badclip ; liang2024poisoned ; liang2024vl ; zhu2024breaking , and jailbreaks. Among them, Jailbreak attacks refer to an attacker’s ability to bypass a model’s built-in safety mechanisms by intentionally designing input prompts or interactions that induce the generative model to produce harmful or illegal content dang2024diffzoo ; deng2023divide ; dong2024jailbreaking ; yang2024mma ; yang2024sneakyprompt ; ying2024jailbreak ; YingSafe2024 ; WangTrojan2025 . The earliest jailbreak attacks emerged in the domain of large language models (LLMs), where attackers employed white-box approaches based on gradient or logits analysis and fine-tuning, as well as black-box approaches such as template injection, prompt rewriting, or leveraging LLMs’ own generations to craft auxiliary inputs, thereby constructing adversarial prompts that elicit prohibited outputs. In the context of T2V, T2VSafetyBench miao2024t2vsafetybench introduces a malicious prompt benchmark covering 12 security dimensions, systematically exposing the widespread vulnerabilities across current commercial and open-source video generation models. SafeWatch chen2024safewatch similarly proposes a benchmark focused on 8 security dimensions and additionally collects a large corpus of prompts containing potentially offensive keywords across 8 novel task categories, enabling indirect assessment of security risks in generated videos. Compared to static text or images, video, as a high-dimensional modality that integrates both visual and temporal dynamics, exhibits significantly greater dissemination power and psychological impact for harmful content. Therefore, jailbreak risks in T2V models are widely recognized as a critical and emerging security challenge for generative AI.
The primary objective of jailbreak defense is to prevent adversaries from successfully inducing models to produce unsafe content xu2024comprehensive ; yi2024jailbreak ; zeng2024autodefense ; xiong2024defensive . Existing research on jailbreak defenses predominantly focuses on LLMs and typically organizes defense mechanisms according to the model processing pipeline into three phases: input stage, model stage, and output stage. At the input stage, defenses include prompt detection, prompt perturbation, and system prompt enforcement. At the model stage, researchers attempt to mitigate jailbreak risks through supervised fine-tuning, adversarial training, and internal model behavior analysis. At the output stage, constructing specialized classifiers or utilizing LLM-based output auditing mechanisms has become a prevalent strategy.
To address the current lack of systematic jailbreak defenses in the T2V domain, this paper proposes a comprehensive and model-agnostic defense framework designed to safeguard T2V models against jailbreak attacks. It examines a variety of feasible defense strategies spanning the three critical stages of T2V security (i.e., input, model, and output stages) and reveals inherent limitations in existing defenses. This framework can also be extended to integrated world simulator systems that generate audio-visual outputs. Most relevant to this work is the I2T (Image-to-Text) jailbreak defense method SAFREE yoon2024safree , which, although evaluated on some T2V models, exhibits limitations due to its reliance on model architecture and parameter access, making it less applicable to closed-source commercial systems.
3 Threat Model
3.1 Victim’s Model
The T2V model aims to learn a generative mapping $G$ from natural language descriptions $\bm{x}\in\mathcal{X}$ to dynamic visual content $\bm{v}\in\mathcal{V}$. The input text space is denoted as $\mathcal{X}$ and the output video space as $\mathcal{V}$. The process of generation by the T2V model can be expressed as $\bm{v}=G(\bm{x})$. The generated video sequence can be denoted as
$\bm{v}=\{v_{1},v_{2},\ldots,v_{T}\},$ (1)
where the generated video sequence consists of $T$ frames, and each frame $v_{t}$ is a static image with dimensions $H\times W\times C$, where $H$, $W$, and $C$ denote the height, width, and number of channels (e.g., RGB three channels), respectively.
Overall, the videos generated by the T2V model aim to obey three rules: (1) Semantic plausibility. The generated video should accurately reflect the semantics of the input text description. (2) Visual quality. The generated frames should have the appropriate visual fidelity. (3) Temporal coherence. The motion between the contents of consecutive frames should be smooth and physically reasonable to ensure dynamic coherence over time.
3.2 Attacker’s Goal
A jailbreak attack is to induce the T2V model $G$ to generate anomalous frames or sequences of frames in the output video $\bm{v}$ that violate the original security constraints through a carefully crafted malicious input text $\bm{x}$. Specifically, the attacker aims for at least one frame $\hat{v}_{t}$ in the generated video $\hat{\bm{v}}$ to satisfy a violation condition. This can be expressed as
$\exists t\in\{1,\dots,T\},\quad\text{R}(\hat{v}_{t})\geq\tau_{R},$ (2)
where $\hat{v}_{t}$ denotes the $t$-th frame in the generated video $\hat{\bm{v}}$. $\text{R}(\cdot)$ is a violation detection function that assesses whether a single frame poses a security risk. $\tau_{R}$ is the predefined risk threshold for violation determination.
In addition, since certain T2V models may implement input filtering mechanisms to block text prompts $\hat{\bm{x}}$ containing sensitive keywords, the attack must also satisfy a stealth constraint. This means that the crafted input text must evade censorship detection and avoid being intercepted. This condition can be expressed as
$\text{D}(\hat{\bm{x}})\leq\tau_{D},$ (3)
where $\text{D}(\cdot)$ is the censorship detection scoring function and $\tau_{D}$ is the detection threshold.
In summary, the attacker wants to craft an input text $x$ that evades censorship detection while ensuring that the resulting generated video $v$ contains at least one frame $f_{t}$ whose violation risk assessment $R(f_{t})$ exceeds the predefined risk threshold $\tau$. The objective of the T2V jailbreak attack can be formalized as:
$\min_{\hat{\bm{x}}}\text{D}(\hat{\bm{x}})\quad\text{s.t.}\quad\text{D}(\hat{\bm{x}})\leq\tau_{D}\quad\text{and}\quad\exists t,\;\text{R}(\hat{v}_{t})\geq\tau_{R}.$ (4)
3.3 Attacker’s Capability
For T2V jailbreak attacks, the attacker’s capabilities can be characterized as: (1) Black-box attack assumption. In practical attack scenarios, the attacker has no knowledge of the internal structure, parameter details, or filtering mechanisms of the victim model. This implies that the generation of malicious input texts does not rely on gradients, model weights, or any other internal information of the target system. (2) Input text manipulation. The attacker has unrestricted access to craft input texts for the T2V jailbreak attempt. During the attack process, the attacker can iteratively modify and refine the input text based on observed outputs to increase the likelihood of successfully inducing the generation of harmful video content. (3) Output observation. The attacker is able to observe whether the victim model generates malicious or non-compliant video outputs, as well as assess the quality of the generated videos. This observation enables the attacker to adjust and optimize subsequent attack inputs based on feedback from the model’s outputs.
4 Defense Framework
4.1 Overview
Inspired by the classification framework of LLM jailbreak defenses, we explore potential defense strategies based on the runtime flow of the T2V model across three dimensions: input, model layer, and output stages (as illustrated in Fig. 2).
In this paper, we mainly evaluate and compare nine defense strategies from the above three domains. Specifically, we investigate five key defense attributes, including (1) commercial deployment adaptability: whether the defense method can be applied to closed-source or black-box models; (2) real-time applicability: whether the defense can be deployed in real-time during the inference phase without introducing unacceptable latency; (3) model access requirement: whether the defense method requires access to, or modification of, the internal model architecture; (4) training dependency: whether the defense relies on additional training or fine-tuning; (5) inference overhead: the additional computational burden introduced by the defense method. As shown in Tab. 1, significant differences are observed among the various defense strategies. In particular, input-level and output-level defense methods generally demonstrate better compatibility with closed-source systems and stronger real-time applicability. Model-side defense methods, while possessing greater theoretical potential, typically require access to internal model structures or retraining, making them impractical for direct application to closed-source T2V systems.
Table 1: Deployment attributes of defense methods in the T2VShield framework.
Category Method Commercial Deployment Inference Overhead Model Access Required Training Stage Required Real-time Applicability
Keyword Filtering $\checkmark$ Low $\times$ $\times$ $\checkmark$
Input-level Implicit Meaning Analysis $\checkmark$ Medium $\times$ $\times$ $\checkmark$
Sensitive Word Segmentation $\checkmark$ Low $\times$ $\times$ $\checkmark$
Adversarial Training $\times$ - $\checkmark$ $\checkmark$ $\times$
Model-level Forgetting Learning $\times$ - $\checkmark$ $\checkmark$ $\times$
SAFREE $\times$ Medium $\checkmark$ $\times$ $\checkmark$
Video Classifier Detection $\checkmark$ Medium $\times$ $\checkmark$ $\checkmark$
Output-level Text Ambiguity Detection $\checkmark$ Low $\times$ $\checkmark$ $\checkmark$
GPT-4o Assisted Review $\checkmark$ High $\times$ $\times$ $\checkmark$
Building on these, the subsequent sections elaborate on the design principles and empirical defense effectiveness of the proposed defense strategies, structured across the three stages.
4.2 Input-level Defense
Input-level defense aims to reduce the potential harm of input prompts or even prevent harmful prompts from being processed by proactively detecting, interpreting, or rewriting the input. We design corresponding defense mechanisms from three perspectives: keyword detection, implicit meaning analysis, and sensitive word segmentation.
Keyword detection. To identify potentially insecure or sensitive content within the input prompt, this defense employs a sensitive content detection mechanism based on keyword matching. Specifically, we construct a collection of sensitive keywords, denoted as $\mathcal{K}$, based on the sensitive keyword lexicon yang2024sneakyprompt . For each input text $\bm{x}$, we parse its content line-by-line and check for the presence of any matching sensitive keywords. The sensitivity determination condition can be formulated as
$\exists k\in\mathcal{K},\quad k\subseteq\bm{x}.$ (5)
That means if there exists any keyword $k$ in $\mathcal{K}$ that is a substring of the input text $\bm{x}$, the text is labeled as sensitive and blocked from proceeding to subsequent processing stages.
Implicit meaning analysis. To gain a deeper understanding of potential hidden harmful content in the text, this defense introduces a harmfulness detection mechanism based on a pre-trained language model. Specifically, we input the text $\bm{x}$ into Toxic-BERT Detoxify , a harmfulness detection model, and use its output harmfulness score as the value of the implicit meaning detector $H(\bm{x})$. The output of the detector $H(\bm{x})$ represents the predicted probability that the input text $\bm{x}$ belongs to the “toxic” category, ranging from $[0,1]$. We set the harmfulness determination threshold as $\tau_{H}$ (default value as 0.5), and the implicit harmfulness detection condition can be formalized as
$H(\bm{x})\geq\tau_{H}.$ (6)
When the output score $H$ is greater than or equal to the threshold $\tau_{H}$, the input text $\bm{x}$ is considered to exhibit significant implied harmful content and is subsequently blocked.
Sensitive word segmentation. This defense module introduces a sensitive word segmentation mechanism to reduce potential input toxicity by rewriting the input prompt. Specifically, given the input text $\bm{x}$ and the set of sensitive words $\mathcal{K}$, we first check whether there exists any sensitive word $k\in\mathcal{K}$ that appears in $\bm{x}$. If a match is found, we apply a syncopation operator $\sigma(\cdot)$ to replace the sensitive word $k$ with its segmented form as
$\bm{x}^{\prime}=\sigma(\bm{x}),$ (7)
where $\sigma(\bm{x})$ denotes the application of special segmentation symbols (e.g., “-”, “.”, “*”) to the sensitive word. Finally, the segmented input $\bm{x}^{\prime}$ is fed into the model for subsequent inference.
4.3 Model-level Defense
Adversarial training. This approach enables models to learn to identify and reject potentially malicious inputs by incorporating adversarial prompts with attack intents into the training process. We propose an adversarial training mechanism tailored for the T2V task, which guides the model to produce predefined warning videos upon detecting toxic cue words, instead of generating harmful content. More training details are provided in Sec. 1 of the Supplementary Materials.
Specifically, the training dataset comprises two types of samples: normal samples $(\bm{x},\bm{v})$, where $\bm{x}$ denotes a benign input prompt and $\bm{v}$ is the corresponding normal video; and adversarial samples $(\bm{x}^{\prime},\bm{v}^{\text{safe}})$, where $\bm{x}^{\prime}$ represents a toxic prompt and $\bm{v}^{\text{safe}}$ is a predefined safety video, such as a warning screen displaying the message “This is an unsafe video.” The clean video data are sourced from the publicly available OpenVid-1M dataset, which comprises 300 pairs of randomly selected prompts and their corresponding videos. The malicious prompt data are derived from T2VSafetyBench (280 prompts randomly sampled) and the official SAFREE dataset (20 prompts randomly sampled). The associated harmful videos are generated using the CogVideoX model. The initial adversarial prompts are constructed from a set of unsafe conceptual terms $\mathcal{T}_{\text{unsafe}}$, which includes keywords related to pornography, violence, politically sensitive content, and other restricted topics. To further strengthen the attack potential of these prompts, we adopt a BERT-based adversarial word-level perturbation strategy. Specifically, given an initial adversarial prompt $\bm{x}^{\prime}=[w_{1},w_{2},\dots,w_{k}]$, we randomly select a word $w_{t}$ at each iteration and replace it with a word $\tilde{w}_{t}$ from $\mathcal{T}_{\text{unsafe}}$, resulting in a candidate prompt $\tilde{\bm{x}}^{\prime}$. If the loss function $\mathcal{L}(G(\tilde{\bm{x}}^{\prime};\bm{\bm{\theta}}),\bm{v}^{\text{safe}})$ of the substitution is smaller than that of the original prompt, the substitution is considered more effective and retained.
We denote the T2V generative model as $G$, with parameters $\bm{\bm{\theta}}$. The adversarial training objective is formulated as a minimax problem:
[…]
$+$ $\mathbb{E}_{(\bm{x}^{\prime},\bm{v}^{\text{safe}})}\left[\max_{\tilde{\bm{x}}^{\prime}\in\mathcal{A}(\bm{x}^{\prime})}\mathcal{L}(G(\tilde{\bm{x}}^{\prime};\bm{\bm{\theta}}),\bm{v}^{\text{safe}})\right]\},$
where $\mathcal{A}(\bm{x}^{\prime})$ denotes the set of adversarial variants of the toxic prompt $\bm{x}^{\prime}$, generated under a predefined perturbation budget (e.g., single-word substitution). This formulation ensures that the model learns to correctly respond to the strongest adversarial prompts within this neighborhood, thereby enhancing its robustness against jailbreak attempts.
Unlearning defense. This technique aims to guide the model towards forgetting high-risk content. Specifically, we focus on targeting specific harmful content for loss augmentation, enabling the model to gradually forget the ability to generate such content while preserving its ability to generate normal content. Our clean video-text pairs $(\bm{x},\bm{v})$ and adversarial video-text pairs $(\bm{x}^{\prime},\bm{v}^{\prime})$ are aligned with the adversarial training dataset. More training details are provided in Sec. 1 of the Supplementary Materials.
For normal videos, the optimization objective is to minimize the loss in order to maintain the quality of normal content generation. For malicious videos, the loss is maximized, thereby encouraging the model to forget its ability to generate such content during training. The overall training objective is shown as
$\min_{\bm{\theta}}$$\Big{\{}\mathbb{E}_{(\bm{x},\bm{v})}\left[\mathcal{L}(G(\bm{x};\bm{\theta}),\bm{v})\right]$ (9)
$-\mathbb{E}_{(\bm{x}^{\prime},\bm{v}^{\prime})}\left[\mathcal{L}(G(\bm{x}^{\prime};\bm{\theta}),\bm{v}^{\prime})\right]\Big{\}},$
where $\bm{x}^{\prime}$ denotes harmful inputs containing high-risk cue words, and $\bm{v}^{\prime}$ denotes the corresponding malicious video.
Training-free defense. This type of defense is designed to resist jailbreak attacks without the need for retraining. SAFREE yoon2024safree is a training-free defense method extended to T2V tasks that preserves the original model parameters while mitigating harmful concepts in the generated content through a dual-stage filtering strategy applied in both the textual embedding space and the visual latent space. This dual-space cooperative defense provides a good balance between generation quality and interpretability. However, SAFREE requires deep integration into the text encoder and denoising module of diffusion-based models. This dependency on internal representations limits its applicability to certain model architectures. For instance, in end-to-end Transformer-based systems such as OpenSora, where there is limited access to intermediate representations during generation, it is challenging to directly apply SAFREE’s strategy. As a model-integrated representation correction mechanism, SAFREE depends on prior knowledge of the model structure and is inherently constrained by architectural compatibility.
4.4 Output-level Defense
The goal of this defense is to establish a multi-dimensional censorship mechanism to achieve accurate identification and real-time blocking of offending content through systematic content analysis and risk detection of generated video sequences. Aiming at the high-dimensional characteristics and information complexity of video data, we intend to explore the feasibility of detection at the level of visual classification, visual semantics, and a comprehensive review by the open-source model.
Visual classification. To automate the identification and annotation of potentially hazardous content in generated videos, we adopt a lightweight video content analysis strategy based on an image classification model. The method evaluates the overall security of a video by randomly sampling video frames and utilizing the open-source insecure content recognition model  man to perform frame-by-frame classification on static frame images.
Specifically, given a video to be analyzed $\bm{v}$, we evenly sample $N=10$ frame images from it:
$\{v_{1},v_{2},\dots,v_{N}\}\sim\text{SampleFrames}(\bm{v}).$ (10)
Input each frame into the well-trained MobileNet v2 model $f(\cdot)$ to obtain its classification label $y_{i}\in\{\text{safe},\text{unsafe}\}$:
$y_{t}=f(v_{t}),\quad t=1,2,\dots,N.$ (11)
The final video’s safety is determined by the following rule:
$\text{Label}(\bm{v})=\begin{cases}\text{unsafe},&\text{if }\exists t,\,y_{t}=\text{unsafe}\\ \text{safe},&\text{otherwise}\end{cases}$ (12)
Video semantic detection. Benefiting from the development in the field of multimodal video understanding, we adopt a semantic-level video comprehension approach to transform video frame sequences into natural language descriptions in order to enhance the detection of unsafe content. Compared to traditional single-frame classification methods, such strategies can effectively reduce the false positive rate and synthesize the high-dimensional semantic information of videos in the temporal dimension.
Specifically, given a generated video $\bm{v}$, we first use the BLIP-2 model  li2023blip to extract the visual features of several of its key frames $v_{1},v_{2},\dots,v_{N}\}$ and feed them into the instruction-tuned MA-LMM model he2024ma , to generate the semantic description text of the video $\bm{c}=\text{Caption}(\bm{v})$. Subsequently, we use the pre-trained natural language inference model BART-large-MNLI lewis2019bart to perform semantic risk assessment on this description text $\bm{c}$ and output the discriminative scores in two dimensions: (1) category confidence vector: $\bm{p}=(p_{\text{safe}},p_{\text{porn}},p_{\text{violence}},\dots)$, indicating the probability that the semantics belongs to each risk category; and (2) ambiguity score: $\alpha(\bm{c})\in[0,1]$ measuring the presence of evasive or vague expressions in the text (e.g., the use of ambiguous or implicit language to convey violations). The decision rule for the final video is shown as follows:
$\text{Label}(\bm{v})=\begin{cases}\text{unsafe},&\text{if }\max\limits_{k\in\mathcal{C}_{\text{risky}}}p_{k}>0.7\\ \text{potential unsafe},&\text{if }\alpha(\bm{c})>0.7\\ \text{safe},&\text{otherwise}\end{cases}$ (13)
where $\mathcal{C}_{\text{risky}}$ denotes the set of high-risk categories (e.g., pornography, violence, gore). Here, we consider both unsafe videos and potentially unsafe videos as malicious outputs.
Model checking. In recent years, automated risk assessment of generated content using large language models (e.g., GPT-4o, Qwen) has gradually become an important defense paradigm. With these powerful general-purpose models, we can achieve semantic-level moderation of potentially harmful content without relying on additional training or manual annotation. Here, we adopt the evaluation protocol of T2VSafetyBench miao2024t2vsafetybench and utilize GPT-4o to evaluate the security of the generated video $\bm{v}$. Specifically, we first extract image frames $\{\bm{v}_{1},\bm{v}_{2},\dots,\bm{v}_{T}\}$ from the video at a frequency of 1 frame per second, and combine them with the preset system prompts to construct the input context $\bm{x}_{\text{LLM}}$ for processing by GPT-4o. The model will give a score $\rho(\bm{v})\in[0,1]$ on whether it contains unsafe information based on the overall video semantics and image content. The risk judgment criteria for the final video are:
$\text{Label}(\bm{v})=\begin{cases}\text{unsafe},&\text{if }\rho(\bm{v})\geq 0.6\\ \text{safe},&\text{otherwise}\end{cases}$ (14)
where $\rho(\bm{v})$ denotes the unsafe score returned by GPT-4o. If a video is determined as “malicious video”, the system will refuse to display or propagate the video content downstream, thus effectively mitigating the negative consequences of jailbreak attacks.
4.5 Preliminary Experiments and Key Insights
Table 2: Performance comparison of different jailbreak defense methods against the CogVideoX model on T2vSafebench.
Method Jailbreak Clean
lr)5-7 ASR(↓) GPT-4o Score(↓) Human Score(↓) CLIP Similarity (↑) Temporal Consistency (↑) FVD Distance (↓)
Undefended 55.6% 52.6 47.7% 0.299 0.966 474.6
Keyword Detection 44.1% 41.6 36.7% 0.294 0.965 473.5
Implicit Meaning Analysis 54.5% 52.4 45.5% 0.289 0.963 48.5
Sensitive Word Segmentation 47.3% 45.3 40.9% 0.297 0.978 482.3
Adversarial Training 0.0% 0.0 0.0% 0.145 0.975 1268.0
Unlearning Defense 0.1% 0.7 0.0% 0.157 0.969 1469.5
Training-free Defense 55.3% 53.2 46.5% 0.293 0.966 474.6
Visual Classification 54.9% 52.0 45.3% 0.290 0.963 485.2
Video Semantic Detection 53.9% 50.6 46.6% 0.296 0.965 476.1
Model Checking 47.2% 44.3 40.3% 0.298 0.963 473.2
Preliminary experiments setting. To preliminarily assess the performance differences among the three types of defense mechanisms in terms of robustness and utility, we select a subset of the T2VSafetyBench jailbreak dataset on the CogVideoX model for experimental evaluation. Specifically, we randomly sample 20 adversarial prompts from each risk category in this dataset, yielding a total of 280 samples to test the effectiveness of the defense against malicious inputs. In addition, to evaluate the impact of each defense type on the model’s generation ability in non-adversarial scenarios, we randomly sample 300 benign text prompts from the MSVD dataset chen2011collecting as clean inputs for measuring video quality.
We adopt six evaluation metrics to measure both the security and generation quality of the T2V system. The security-related metrics include: (1) Attack Success Rate (ASR); (2) GPT-4o Automated Risk Assessment Score; (3) Human Evaluation (Human). Lower values on these three metrics indicate better defense effectiveness and a lower likelihood of generating harmful content. For video quality evaluation, we consider three aspects: (4) CLIP Similarity, which measures the semantic alignment between the video and the input text; (5) Temporal Consistency, which reflects the coherence of frame transitions over time; (6) Fréchet Video Distance (FVD), which quantifies the distributional distance between generated and real videos. Notably, a lower FVD score indicates higher perceptual similarity to real videos. The formal definitions and computational details of the six metrics are provided Sec. 6.1.
Defense results and insights. Tab. 2 shows the performance differences of nine representative T2V defense methods on the jailbreak subset (T2VSafetyBench) and the clean subset (MSVD). From the results, we can draw the following key observations and conclusions.
❶ Although there exist recognizable risk signals within the input prompts, the effectiveness of current input-layer defenses remains limited. For instance, among the three input-side approaches, Keyword Detection lowers the ASR from 55.5% to 44.1% (an 11.4% reduction), yet still permits a substantial number of successful jailbreaks. Similarly, Sensitive Word Segmentation and Implicit Meaning Analysis only reduce the ASR to 47.3% and 54.5% respectively. This suggests that although structured risk cues are embedded in prompts, conventional detection methods fail to effectively extract and generalize these patterns, thus only defending against a subset of attack samples. This phenomenon reveals that while current input defenses can capture some explicit threats, they do not yet systematically leverage the structural and semantic patterns embedded within the prompts, which results in rewriting strategies that are insufficient to disrupt the attack pipeline. {abox} Insight 1: Prompt-based vulnerabilities. Jailbreak prompts often contain identifiable semantic cues and structural patterns, yet current defenses fail to systematically utilize them to enhance input robustness.
❷ The temporal sparsity of harmful content limits the efficacy of output-layer frame-based detection strategies. On the output side, while all three methods (Visual Classification, Video Semantic Detection, and Model Checking) slightly reduce ASR and GPT-4o risk scores compared to the undefended baseline (e.g., Model Checking reduces the ASR to 47.2% and the GPT-4o score to 44.3), the overall defense effect remains suboptimal, with human annotation indicating a risk rate above 40.0%. Further analysis shows that these detection methods are often sensitive to isolated anomalies in single frames but struggle to identify dispersed, short-duration yet critical high-risk segments. This suggests that relying solely on uniform-granularity frame-level or semantic-level detection is inadequate for addressing the highly temporally irregular distribution of harmful content in T2V outputs, especially when risks emerge only briefly in isolated moments. {abox} Insight 2: Temporal sparsity and randomness. Harmful content typically manifests in only a few frames or short intervals, exhibiting high sparsity and discontinuity, making it difficult for frame-wise or single-scale detection to maintain stable and high recall.
❸ Model-centric defenses, although theoretically effective, often degrade generation quality and lack practicality. For example, both Adversarial Training and Unlearning Defense reduce the ASR to 0%, but their FVD scores reach 1268.0 and 1469.5 respectively—much higher than that of the original system (474.6). Despite their strong robustness, the high computational cost and intrusive modifications to model architecture severely impact the generation quality under clean inputs, limiting their applicability in real-world deployment. {abox} Insight 3: Model-centric constraints. While training-level defenses offer strong theoretical robustness, their high cost, disruptive nature, and poor compatibility with closed-source models pose major challenges for real-world deployment.
Given that traditional staged defenses still exhibit residual risks and inconsistent performance in empirical evaluations, we propose the T2VShield framework, which provides a unified approach to modeling risks at both the input and output levels. The details of this framework are elaborated in Sec. 5.
5 T2VShield Defense
Figure 2: Overview of the defense framework for mitigating jailbreak attacks in T2V. It integrates three stages of defense: (1) Input-level defense, (2) Model-level defense, and (3) Output-level defense. The proposed T2VShield framework includes the following modules: At the input level, we introduce a large-model-driven reasoning-chain rewriting method (RiskTrace CoT) with positive and negative samples (PosNegRAG) to achieve a more interpretable and reliable rewriting strategy. On the output level, we design a multi-scope output detection mechanism, which jointly evaluates the semantic and content layers of the generated video through multi-timescale segmentation and cross-modal feature integration.
5.1 Motivation and Design
Based on the above inspiration, we recognize the necessity of proposing a model-agnostic jailbreak defense approach to address model-centric constraints. To this end, we shift the focus of our defense design to the input and output stages rather than relying on modifications to the internal structure of the T2V model. By combining two key observations: ❶ the prevalence of parsable semantic cues and structural patterns in existing jailbreak prompts, and ❷ the fact that harmful content in the generated videos tends to exhibit temporal sparsity and transient anomalies, we propose the T2VShield defense, as shown in Fig. 2.
T2VShield systematically defends against jailbreak attacks on T2V models through a two-phase defense strategy. To proactively mitigate potential risks at the input stage, T2VShield adopts a two-pronged approach: (1) it conducts structured risk tracing and targeted rewriting through RiskTrace CoT (Sec. 5.2), utilizing chain-of-thought reasoning to reduce input risk exposure; and (2) it introduces PosNegRAG (Sec. 5.3), a retrieval-augmented method that provides concrete positive and negative examples to enhance rewriting reliability and mitigate the inference instability of multimodal models. For the output stage, T2VShield enhances risk detection by combining multi-timescale slicing and multimodal feature fusion (Sec. 5.4), a strategy we term Multi-scope detection, to systematically improve coverage and accuracy across both temporal and semantic levels.
Overall, our defense framework can achieve generalizable protection for T2V models under diverse jailbreak attack scenarios and can be seamlessly applied across different T2V systems.
5.2 Rewriting Prompt via RiskTrace CoT
Jailbreak attackers often embed implicit guiding signals into the input prompts in the form of semantic hints and structural patterns, aiming to induce the T2V model to generate harmful content. We have revealed that jailbreak prompts have parsable weaknesses at the semantic and structural levels in Insight 1. However, existing defense methods mostly focus on filtering sensitive words and shallow rewriting at the input level, failing to systematically exploit these structural cues and resulting in limited defense effectiveness.
Therefore, we propose a chain-of-thought reasoning rewriting strategy based on risk tracing (RiskTrack CoT). The entire process is performed by a multimodal large language model $\mathcal{M}$ and consists of four stages: (1) Risk comprehension via prompting. A surface-meaning and intent understanding prompt is issued to $\mathcal{M}$, prompting it to generate a reasoning chain $\bm{r}=\mathcal{M}_{\text{reason}}(\bm{x})$ that captures the explicit meaning and inferred risk intent underlying the input prompt $\bm{x}$. (2) Risk element identification via prompting. A risk element extraction prompt is then applied to $\mathcal{M}$ based on the reasoning chain $\bm{r}$, resulting in the identification of potential risk elements $\mathcal{E}$ and their corresponding mitigation strategies $\mathcal{S}$, formalized as $(\mathcal{E},\mathcal{S})=\mathcal{M}_{\text{identify}}(\bm{r})$. (3) Directed rewriting via prompting. Using a targeted rewriting prompt, $\mathcal{M}$ rewrites the input $\bm{x}$ into a safer version $\bm{x}^{\prime}=\mathcal{M}_{\text{rewrite}}(\bm{x},\mathcal{E},\mathcal{S})$ by removing, replacing, or abstracting the identified risks while preserving the benign semantics. (4) Safety verification via prompting. Finally, a self-check prompt is issued to evaluate whether $\bm{x}^{\prime}$ still contains residual risks. Only rewritten prompts passing this final validation are retained.
By breaking down the defense into an explicit multi-stage prompting process, RiskTrace CoT enables a fine-grained, interpretable, and robust prompt rewriting mechanism to defend against jailbreak attacks. More detailed examples of prompt templates and reasoning chains are provided in the supplementary materials.
5.3 Assisted Rewriting via PosNegRAG
Although current advanced multimodal large language models (e.g., GPT-4o, LLaVA) possess basic risk identification and rewriting capabilities, we observe significant performance discrepancies in real defense tasks, which manifest in two key issues: (1) Principle adherence deficiency. Models struggle to accurately understand defense objectives when only abstract safety instructions are provided without concrete examples, leading to ineffective rewriting. (2) Capability ceiling effect. Due to differences in size and training, models exhibit limits in complex reasoning and fine-grained risk perception, making CoT-based rewriting success heavily model-dependent. To alleviate these limitations, we propose PosNegRAG, a retrieval-augmented rewriting strategy that supplements the rewriting model with concrete positive and negative examples as guidance. Next we will describe the construction and retrieval of PosNegRAG.
Sample pool construction. We construct a positive-negative sample pool based on the SafeWatch dataset chen2024safewatch , selecting 143 clean video-text pairs as positive examples and 717 harmful ones as negative examples. Each sample $\bm{s}_{i}$ is represented by a joint feature pair $(\bm{z}_{i}^{\text{text}},\bm{z}_{i}^{\text{image}})$, where $\bm{z}_{i}^{\text{text}}=E_{\text{text}}(\bm{s}_{i}^{\text{text}})$ is obtained using SentenceTransformer, and $\bm{z}_{i}^{\text{image}}$ is the average-pooled CLIP encoding over $n=4$ sampled frames:
$\bm{z}_{i}^{\text{image}}=\frac{1}{n}\sum_{k=1}^{n}E_{\text{image}}(\bm{s}_{i}^{\text{image},(k)}).$ (15)
Multimodal retrieval graph construction. Samples are treated as nodes, and pairwise multimodal similarity is computed as:
$\text{sim}(\bm{s}_{i},\bm{s}_{j})=$ $\alpha\times\cos(\bm{z}_{i}^{\text{text}},\bm{z}_{j}^{\text{text}})$ (16)
$+(1-\alpha)\times\cos(\bm{z}_{i}^{\text{image}},\bm{z}_{j}^{\text{image}}),$
where $\alpha=0.7$ balances text and image features.
Edges are established based on similarity thresholds:
$\mathcal{E}_{\text{intra}}=\{(\bm{s}_{i},\bm{s}_{j})\mid y_{i}=y_{j},\ \text{sim}(\bm{s}_{i},\bm{s}_{j})>\tau_{\text{pos}}\},$ (17)
$\mathcal{E}_{\text{inter}}=\{(\bm{s}_{i},\bm{s}_{j})\mid y_{i}\neq y_{j},\ \text{sim}(\bm{s}_{i},\bm{s}_{j})>\tau_{\text{neg}}\},$ (18)
where $y_{i}$ and $y_{j}$ denote the class labels (positive or negative) of nodes $\bm{s}_{i}$ and $\bm{s}_{j}$, respectively. The intra-class threshold $\tau_{\text{pos}}=0.7$ and inter-class threshold $\tau_{\text{neg}}=0.3$. The final multimodal retrieval graph is denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{E}=\mathcal{E}_{\text{intra}}\cup\mathcal{E}_{\text{inter}}$.
Example retrieval and integration. Given an input prompt $\bm{x}$, we first encode its text feature $\bm{z}_{x}=E_{\text{text}}(\bm{x})$. Negative examples are retrieved based on a combined score incorporating direct similarity and graph-based connectivity as
$\text{score}(\bm{s}_{j})=$ $\cos(\bm{z}_{x},\bm{z}_{j}^{\text{text}})+\lambda\times\frac{1}{\mid\mathcal{N}\mid}$ (19)
$\sum_{\bm{s}_{k}\in\mathcal{N}}\mathbb{I}\left[(\bm{s}_{j},\bm{s}_{k})\in\mathcal{E}_{\text{intra}}\right]\times\text{sim}(\bm{s}_{j},\bm{s}_{k}),$
where $\lambda=0.2$ controls the weight of graph connectivity. Subsequently, for each retrieved negative sample, its closest positive neighbor is selected along the strongest inter-class edge as
$\mathcal{N}_{\text{pos}}(\bm{x})=\Big{\{}\bm{s}_{k}\mid\bm{s}_{k}=\arg\max_{(\bm{s}_{j},\bm{s}_{k}^{\prime})\in\mathcal{E}_{\text{inter}}}\text{sim}(\bm{s}_{j},\bm{s}_{k}^{\prime})\Big{\}}.$ (20)
Finally, the auxiliary positive and negative examples $(\mathcal{N}_{\text{pos}}(\bm{x}),\mathcal{N}_{\text{neg}}(\bm{x}))$ are injected into the rewriting process to guide safer prompt generation.
5.4 Multi-scope Output Detection
We observe in Insight 2 that harmful content often exhibits temporal sparsity and transient anomaly characteristics in T2V outputs and is exposed only in a few frames or short temporal segments. This characteristic poses two challenges. First, a single time-scale or uniform frame sampling strategy is insufficient to cover all risky regions, resulting in inadequate detection coverage; second, a single-modality detection approach is unable to capture both local details and global semantics, further weakening detection accuracy. To address these challenges, we propose Multi-scope Output Detection, which improves detection coverage and accuracy through multi-timescale slicing and cross-modal feature fusion, thereby enhancing the overall stability of output-side risk detection.
Multi-timescale slicing. Considering that harmful content typically exhibits sparse distribution and transient anomalies in the temporal dimension, a single-scale frame sampling strategy can easily lead to undetected risky content. To this end, we introduce a multi-timescale slicing mechanism to comprehensively model the content dynamics of video $\bm{v}$ at different temporal granularities. Specifically, video $\bm{v}$ is partitioned into three temporal levels: global (full video), mesoscale (approximately 15-frame clips), and fine-grained (approximately 5-frame clips). Within each scale, cross-modal features are extracted for subsequent detection using overlapping segmentation with uniform frame sampling. This multi-granularity temporal modeling approach balances the sensitive capture of local anomalies with the holistic perception of long-term dynamics, thereby significantly improving detection coverage.
Multimodal feature fusion. Compared to local image-based detection, text semantics can summarize the overall content of the video at a higher abstraction level, helping to identify potential risks that are difficult to capture through local visual analysis alone. To further enhance detection accuracy, we synchronize the extraction of information from both image and text modalities to build a complementary detection mechanism. Specifically, for the sampled set of image frames $\{v_{1},\dots,v_{n}\}$, we evaluate the content of each frame using an image-level NSFW detector to capture local visual anomalies. Simultaneously, the image frames are input into a multimodal vision-language model (e.g., BLIP-2 Vicuna) to generate corresponding natural language descriptions, and the overall semantic-level potential risk is assessed through text ambiguity detection with an offending content-level classifier. Within each segment, if any modality detects a violation, the segment is considered risky; at the video level, we adopt a “local-to-global subsumption” strategy, meaning that if any segment is detected as harmful, the entire video $\bm{v}$ is classified as unsafe. This strategy effectively improves the overall detection accuracy and robustness.
Our proposed method possesses strong interpretability and security robustness. Since T2VShield operates solely at the input and output levels, it exhibits cross-model compatibility and robustness, making it broadly applicable to jailbreak defense tasks across various T2V systems. Specific algorithmic details are provided in Sec.2 of the Supplementary Material.
6 Experiment and Evaluation
6.1 Experiments Setting
Datasets. To comprehensively evaluate the performance of the proposed T2VShield framework in terms of jailbreak defense and video quality preservation, we utilize three representative datasets: T2VSafetyBench miao2024t2vsafetybench , SafeWatch chen2024safewatch , and MSVD chen2011collecting , each corresponding to a specific evaluation scenario—jailbreak input detection, general malicious content defense, and clean input quality assessment, respectively. (1) T2VSafetyBench serves as our primary benchmark for jailbreak evaluation. It covers 14 high-risk categories, including pornography, violence, political sensitivity, copyright infringement, and temporal risk. We randomly sample 20 adversarial prompts from each category, resulting in 280 total attack cases to simulate representative jailbreak scenarios. (2) SafeWatch is a large-scale video security inspection dataset that includes over 2 million video samples spanning six major risk categories. We randomly select 50 representative malicious text descriptions as an auxiliary test set to assess the generalization and robustness of defense mechanisms across broader risk domains. Note that SafeWatch is not specifically tailored for jailbreak attacks, so some defense methods may trivially succeed on this dataset. (3) MSVD is used to assess the impact of defense strategies on clean input prompts. We randomly select 300 text queries from the MSVD dataset to measure generation quality under benign conditions.
Victim models. We selected two open-source models (Open-Sora and CogVideoX) and three closed-source models (Kling, Luma, and Pika) to evaluate the victimized T2V system. Open-Sora zheg024open is one of the most representative open-source models in the community, which supports the generation of long-duration, high-quality videos, and has good transparency and reproducibility; CogVideoXyang2024cogvideox introduces a 3D spatiotemporal modeling mechanism while maintaining high visual fidelity, producing results with better temporal coherence, which facilitates the analysis of jailbreak attacks on the dynamic structure of the video. In terms of closed-source models, Kling kling2024 emphasizes multi-character storytelling in Chinese-language contexts, Luma luma2024 focuses on the rendering of real-world scenes and lighting effects, and Pika pika2024 is targeted at creative professionals and supports multiple video generation styles. These three closed-source systems have their own strengths in terms of application scenarios, generation styles, and functional design, representing the mainstream trends of commercial T2V applications, and thus offer strong representativeness.
Auxiliary models. We introduce multiple auxiliary models to support key aspects of the defense process. First, to evaluate the safety of the generated videos, we employ GPT-4o to determine the jailbreak success rate and assign a risk score to the video content. In the input defense phase of T2VShield, we primarily use the LLaVA model li2024llava to perform Chain-of-Thought (CoT) reasoning with negative sample-based semantic rewriting; in addition, we further analyze the influence of models such as Qwen yang2024qwen2 , GPT-4o, and DeepSeek liu2024deepseek on rewriting performance in the discussion section. In the output detection phase, MA-LLM he2024ma is used to understand video semantics and produce textual descriptions, while the NSFW detector screens image frames for inappropriate content at the visual level. Finally, when constructing the multimodal retrieval graph and sample representations, we also utilize CLIP radford2021learning to extract visual features to assist in accurate positive-negative sample matching and semantic guidance.
Evaluation Metrics. We use six metrics to evaluate the safety and video quality of the T2V generation system. The safety metrics include: (1) Attack Success Rate (ASR) niu2024jailbreaking : the proportion of generated videos that still produce harmful content despite receiving adversarial prompts. This metric assesses whether the defense mechanism effectively prevents jailbreak attacks. (2) GPT-4o Score miao2024t2vsafetybench : This score is obtained by prompting GPT-4o to conduct multiple rounds of evaluation on the generated videos, outputting a risk level (e.g., 0%-100%). It simulates an advanced LLM performing “content moderation” to evaluate potential violations in the output. A lower value indicates a safer video with less exposure to harmful content. (3) Human-perception Labeling (Human): intuitive human judgment of the security and appropriateness of the video. A lower score implies better defense performance and is often used to corroborate the reliability of automatic evaluation results.
Video quality metrics include: (1) CLIP Similarity radford2021learning : We apply the CLIP model to compute the multimodal similarity between generated video frames and the original input prompt. This reflects how well the video semantically aligns with the intended text; higher similarity means better alignment with user intent. (2) Temporal Consistency varghese2020unsupervised : evaluates the coherence of the generated video along the time axis. It measures whether the video maintains consistent motion, structure, and semantics, and whether it avoids artifacts like “frame skipping” or “flickering”. A higher score indicates smoother and more natural temporal dynamics. (3) Fréchet Video Distance (FVD) unterthiner2019fvd : calculated based on the Fréchet distance between the feature distributions extracted via the I3D model, this metric measures the distributional gap between generated and real-world videos. A lower FVD means that the generated video is closer to real data in perceptual quality and naturalness.
Additionally, we report the time cost (in seconds) required by each defense approach, covering both training and inference stages.
6.2 Defense Results on Jailbreak and Clean Inputs
Table 3: Evaluation results on Open-Sora across jailbreak and clean prompts.
Method T2vSafebench Jailbreak Clean
ASR (↓) GPT-4o Score (↓) Human (↓) CLIP Similarity (↑) Temporal Consistency (↑) FVD (↓)
Undefended 52.9% 50.7 56.0% 0.158 0.972 639.1
Keyword Detection 42.1% 40.29 46.4% 0.158 0.980 643.2
Implicit Meaning Analysis 52.9% 50.7 55.6% 0.159 0.979 645.9
Sensitive Word Segmentation 52.1% 50.54 54.2% 0.167 0.982 632.3
Adversarial Training 34.3% 33.25 37.6% 0.165 0.986 747.1
Unlearning Defense 45.2% 44.571 48.6% 0.165 0.991 919.4
Visual Classification 43.8% 41.57 46.9% 0.159 0.969 699.6
Video Semantic Detection 48.5% 46.83 51.4% 0.156 0.965 687.5
Model Checking 41.9% 40.25 44.4% 0.156 0.968 635.8
T2VShield 17.2% 15.64 19.6% 0.163 0.989 634.5
Tab. 3 shows the defense performance of T2VShield and the nine mainstream defense methods on Open-Sora against the T2VSafetyBench jailbreak sample and the MSVD clean sample, respectively. The ASR, GPT-4o score and human labeling evaluate safety (lower is better), while CLIP similarity, Temporal Consistency, and FVD evaluate video quality (higher/lower is better). Based on the data in the Tab. 3, we draw the following conclusions: ❶ T2VShield significantly outperforms existing methods on safety metrics. Compared with the undefended model (ASR 52.9%, GPT-4o score 50.7, Human score 56.0%), T2VShield reduces the attack success rate to 17.2%, the GPT-4o score to 15.6, and the human evaluation to only 19.6%, which clearly surpasses all baseline methods and demonstrates strong, robust jailbreak defense capability. ❷ T2VShield maintains a strong balance between safety and video quality without compromising clean sample performance. Despite T2VShield’s enhanced input and output defenses, its CLIP similarity (0.163), temporal consistency (0.989), and FVD (634.5) on clean samples are close to those of the undefended model (0.158 / 0.972 / 639.1) and other lightweight defense strategies, suggesting that it enhances safety while preserving th gneration quality under normal inputs. ❸ Training-based in-model defenses suffer from efficiency and quality degradation. Although Adversarial Training reduces ASR to 34.3%, its FVD increases to 747.1, and Unlearning performs even worse, reaching 919.4, indicating significant quality degradation; in contrast, T2VShield achieves better security with lower cost to quality without altering the model structure.
In summary, T2VShield significantly improves resistance to jailbreak attacks while maintaining video quality, offering a better trade-off between safety and usability than traditional approaches. We present visualization results in Sec. 4 of the Supplementary Material.
6.3 Generalization to Real-World Harmful Prompts
Table 4: Comparison of jailbreak defense methods on the SafeWatch dataset using two open-source T2V models.
Method CogVideoX OpenSora
ASR (↓) GPT-4o Score (↓) Human (↓) ASR (↓) GPT-4o Score (↓) Human (↓)
Undefended 47.8% 46.9 52.0% 49.2% 50.55 54.3%
Keyword Detection 25.4% 26.1 28.0% 31.4% 31.25 35.2%
Sensitive Word Segmentation 29.6% 30.1 33.5% 32.5% 31.92 34.5%
Adversarial Training 5.0% 5.25 7.1% 38.8% 37.75 40.3%
Unlearning Defense 0.0% 0.00 0.0% 31.35% 29.25 35.0%
Training-free Defense 45.1% 45.2 48.1% / / /
Visual Classification 46.7% 45.6 48.5% 42.2% 42.37 45.2%
Video Semantic Detection 46.9% 46.1 50.3% 43.2% 42.35 43.2%
Model Checking 38.2% 36.9 39.6% 36.3% 37.26 39.3%
T2VShield 12.3% 11.1 13.0% 16.8% 14.75 18.2%
Tab. 4 presents a comparison of the effectiveness of various defense approaches on two representative open-source T2V models (CogVideoX and Open-Sora), evaluated using the real-world malicious dataset SafeWatch. We randomly sample 50 harmful descriptions from SafeWatch, covering diverse categories of real-world risks, and assess each method’s actual performance under non-synthetic attack scenarios. We summarize the following key findings and conclusions: ❶ T2VShield achieves the most effective protection under real-world malicious inputs. Compared to the high attack success rate in the Undefended setting (47.8% for CogVideoX and 49.2% for Open-Sora), T2VShield reduces this to 12.3% and 16.8%, respectively. It also ranks best in both GPT-4o safety scores and human annotations. This demonstrates that our framework not only handles synthetic jailbreak prompts well but also generalizes effectively to real-world harmful input cases, showcasing its practical applicability. ❷ Most traditional defenses show improved performance on SafeWatch compared to T2VSafetyBench, indicating better responsiveness to explicit threats. For instance, the attack success rate of Keyword Detection decreases from 42.1% (on T2VSafetyBench) to 25.4% (CogVideoX) and 31.4% (Open-Sora), and Model Checking exhibits a similar pattern. This trend suggests that SafeWatch attacks often involve overtly harmful semantics, which can be easily identified by surface-level strategies. However, these approaches perform significantly worse on T2VSafetyBench, which emphasizes structural obfuscation and semantic manipulation, revealing a lack of generalizability. ❸ T2VShield maintains consistent performance across models, reflecting strong cross-model generalization. Unlike other methods that show large performance variation between CogVideoX and Open-Sora, T2VShield delivers leading results on both models (e.g., Human-labeled violations are 13.0% and 18.2%, respectively), highlighting its architecture-agnostic robustness.
In conclusion, these results further validate the real-world applicability of T2VShield. Compared with baseline approaches, our framework demonstrates superior performance in three aspects: safety effectiveness, model adaptability, and input distribution generalization, making it a suitable candidate for a reliable and universal T2V security solution.
6.4 Efficiency–Effectiveness Trade-off Analysis
Table 5: Comparison of defense methods on CogVideoX using the T2VSafetyBench dataset. The table shows training and inference time (per 100 samples) and the corresponding ASRs.
Method Training Time Inference Time Total Time ASR (↓)
Keyword Detection 0s 0.1s 0.1s 44.1%
Implicit Meaning Analysis 0s 3s 3s 54.5%
Sensitive Word Segmentation 0s 5.3s 5.3s 47.3%
Adversarial Training 43866.2s 0s 43866.2s 0.0%
Unlearning Defense 41366.1s 0s 41366.1s 0.1%
Training-free Defense 0s 54500.4s 54500.4s 55.3%
Visual Classification 0s 327.1s 327.1s 54.9%
Video Semantic Detection 0s 428.9s 428.9s 53.9%
Model Checking 0s 523.6s 523.6s 47.2%
T2VShield 0s 1700.8s 1700.8s 22.4%
Tab. 5 reports the total processing time (including training and inference) and the corresponding ASR of various defense strategies when evaluating 100 samples on the T2VSafetyBench dataset. All experiments are conducted on a single H100 GPU to systematically assess the performance of each approach in terms of the trade-off between efficiency and robustness. We summarize the following key observations: ❶ T2VShield offers a strong balance between security effectiveness and computational cost. Compared to the 44.1% attack success rate in the undefended setting, our method reduces the ASR to 22.4%, while maintaining a total inference time of 1700.8 seconds, making it a viable candidate for deployment. ❷ Training-based defenses achieve very low ASR but incur excessive computational costs, limiting their practicality. For instance, Sensitive Word Segmentation and Adversarial Training reduce the ASR to 0.0% and 0.1%, respectively, but require 43,866.2 and 41,366.1 seconds for training—posing challenges for real-world use cases that demand fast iteration and model adaptability. ❸ Some inference-only defenses do not involve training but introduce high latency. For example, Unlearning Defense takes 54,500.4 seconds, and Training-free Defense takes 327.1 seconds for inference alone. Moreover, the ASR of Unlearning Defense is still as high as 55.3%, indicating that its large time cost does not yield corresponding gains in security.
In conclusion, T2VShield maintains low ASR with moderate inference cost, making it particularly suitable for deployment in scenarios with limited resources or broad generalization requirements.
6.5 Closed-Source Model Evaluation
Table 6: Evaluation results across three video generation backends: Pika, Luma, and Keling. Lower score indicates better jailbreak robustness.
Pika Luma Keling
Method ASR GPT-4o Human ASR GPT-4o Human ASR GPT-4o Human
Undefended 31.3% 28.6 32.3% 28.7% 26.2 28.1% 75.5% 72.4 74.4%
Keyword Detection 23.3% 24.5 23.8% 21.5% 23.5 24.8% 56.6% 58.8 59.6%
Implicit Meaning Analysis 25.6% 27.3 28.9% 23.5% 24.7 25.9% 61.9% 61.1 62.3%
Sensitive Word Segmentation 25.7% 26.9 27.8% 24.9% 25.1 26.6% 65.3% 64.2 65.9%
Visual Classification 29.1% 30.2 32.2% 24.4% 26.3 27.3% 75.3% 73.3 75.2%
Video Semantic Detection 30.2% 31.5 33.7% 26.7% 27.7 29.1% 74.1% 74.5 76.8%
Model Checking 28.4% 29.8 30.3% 22.9% 23.9 25.1% 73.7% 73.2 74.9%
T2VShield 14.7% 12.2 14.2% 14.2% 13.0 14.9% 9.1% 9.2 11.3%
Tab. 6 shows the results of our defense performance evaluation on three closed-source video generation systems (Pika, Luma, and Keling), which are used to measure the generality and robustness of T2VShield and other baseline methods under commercial models. We have the following main observations: ❶ T2VShield achieves the lowest Attack Success Rate (ASR), GPT-4o score, and human evaluation score on all three closed-source models. For example, the ASR drops to 14.7% on Pika, 14.2% on Luma, and even drops to only 9.1% on Keling. This result demonstrates that our defense method remains highly adaptable and effective under proprietary platforms.
❷ In contrast, the performance of other methods on the closed-source models varies significantly, especially on Keling, where the effectiveness of most defense strategies declines sharply. For example, the ASR of traditional methods such as Keyword Detection and Visual Classification on the Keling model reaches up to 56%, indicating that these approaches are difficult to generalize across different model architectures and mechanisms, whereas T2VShield’s end-to-end input/output defense strategy demonstrates stronger model-independence.
❸ Although closed-source models such as Luma and Keling typically implement certain built-in safety mechanisms (e.g., prompt filtering, content moderation, etc.), the table shows that they still face considerable jailbreak risk (e.g., the ASR without defense on Keling reaches 75.5%), which suggests that the intrinsic security measures of existing commercial systems remain limited, and plug-in external defense modules are still essential.
In summary, T2VShield not only performs well in open-source settings, but also exhibits strong transferability and robustness across commercial closed-source models, validating its potential as a universal T2V safety defense module.
6.6 Audio-Visual Consistency under Defense
Table 7: Evaluation of audiovisual consistency and adversarial audio generation.
Alignment T2A Match
Method Multi-choice (↓) Binary (↓) Multi-choice (↓)
(Sync) (Generation)
Undefended 81.2% 89.7% 45.3%
T2VShield 50.1% 43.5% 27.7%
To further validate the potential of our defense strategy in multimodal world modeling, we design an experiment on audio-visual consistency and audio-based adversarial generation, as shown in Tab. 7. The experiment consists of two main components:
First, in the audio-visual consistency evaluation (Alignment), we design both Binary and Multiple-choice tasks, where harmful videos are mixed with benign videos and paired with our curated sets of malicious/benign audio clips. Participants are asked to judge whether the video aligns with the audio. The results show that in the undefended condition, participants can easily determine the correspondence between audio and video (achieving up to 89.7% / 81.2% accuracy), whereas under defense, the judgment accuracy drops significantly (43.5% / 50.1%). This demonstrates that the coherence between the output video and its original malicious audio is disrupted after our defense, thereby validating that visual-level defenses can substantially diminish the effectiveness of joint audio-visual attacks.
Second, in the adversarial audio generation matching (T2A Match), we use text-to-audio generation models such as MusicLDM chen2024musicldm to synthesize audio from malicious prompts, and pair these audio clips with either benign or defended videos in a multiple-choice matching task. The experimental results show that although participants can still identify some mismatches, the overall matching accuracy remains low (only 27.7% under T2VShield), substantially lower than in the undefended setting (45.3%). This further suggests that our visual-layer defenses are effective in mitigating audio-prompted generation threats. Additionally, it implies that current audio-video generation models still fall short in achieving tightly aligned multimodal synthesis.
In summary, our T2VShield not only effectively mitigates visual-only jailbreak attempts, but also significantly reduces the success rate of cross-modal audio-visual attacks, laying the groundwork for future research in secure multimodal world modeling. It is worth emphasizing that our method does not require access to model internals and can be seamlessly applied to a wide range of T2V systems, ensuring strong generalization and deployment flexibility. More questionnaire details can be found in Sec. 3 of the Supplementary Materials.
6.7 Ablation Study
Table 8: Ablation of T2VShield components on Open-Sora under T2VSafetyBench jailbreak attacks.
T2vSafebench Jailbreak
Method ASR (↓) GPT-4o Score (↓) Human (↓)
LLaVA 51.1 48.4 49.2
LLaVA + Multiscope 42.2 39.2 40.3
LLaVA + RiskTraceCoT 44.7 41.3 44.8
LLaVA + RiskTraceCoT + PosNegRAG 29.7 28.1 29.3
T2VShield 22.4 18.3 20.24
Ablation study of core components. According to the principal component ablation experiment results in Tab. 8, we gradually added and evaluated the core components of T2VShield on the Open-Sora model of T2VSafetyBench and came to the following conclusions: ❶ Multi-scope output detection significantly improves protection: After adding the Multi-scope output detection module on the basis of using only the basic LLaVA model, the ASR is reduced from 51.1% to 42.2%, and the GPT-4o score and manual score are also significantly reduced, indicating that output-side detection plays a key role in lowering the jailbreak success rate. ❷ RiskTrace CoT input rewriting has a complementary role in identifying potential risky prompts: The introduction of the RiskTrace CoT module alone can reduce the ASR to 44.7%, which is slightly less effective than Multi-scope, indicating that although semantic regulation at the input stage is helpful for defense, it still has certain limitations when used in isolation. ❸ PosNegRAG effectively enhances the stability and reliability of the rewriting model. After combining positive and negative sample retrieval on the basis of CoT, the ASR dropped to 29.7%, further validating the important role of positive and negative examples in enhancing the risk assessment and rewriting robustness of multimodal large models. ❹ T2VShield achieves optimal performance when all modules are integrated: After fully utilizing the RiskTrace CoT, PosNegRAG, and Multi-scope detection modules, the final ASR dropped to 22.4%, and both the GPT-4o and manual evaluations also declined notably, indicating that the three modules work for input-side semantic filtering and output-side multi-scale detection, which is the key to achieving optimal protection. Based on the above results, T2VShield adopts a dual-phase input-output structural design, where each module contributes independently and complementarily to security enhancement, validating the effectiveness and necessity of our approach.
Effectiveness of PosNegRAG.
Table 9: Comparison of different RAG strategies under RiskTraceCoT on the Open-Sora model with the T2VSafetyBench jailbreak dataset.
Jailbreak
Method ASR (↓) GPT-4o Score (↓) Human (↓)
LLaVA 51.1 48.4 49.2
LLaVA + RiskTraceCoT + NegPosRetri 31.4 30.2 33.4
LLaVA + RiskTraceCoT + NegRAG 28.2 25.6 27.8
LLaVA + RiskTraceCoT + NegPosRAG 26.3 22.3 23.7
In Tab. 9, we systematically evaluate the defense effectiveness of the graphical RAG (Retrieval-Augmented Generation) mechanism built on positive and negative samples. We compare three different example retrieval strategies on the T2VSafetyBench dataset using the Open-Sora model as an attack target: (1) NegPosRetri: filtering positive and negative samples by similarity scores only. Considering the worst-case scenario, we assume the input prompt to be malicious and thus use the sample with the highest similarity as a negative example and the lowest as a positive example. (2) NegRAG: A graph structure constructed based on negative samples only, where examples are built based on intra-class similarity, is used to aid negative risk recognition in the rewrite phase. (3) NegPosRAG (our proposed method): Combining both positive and negative samples during the construction process and encoding the semantic relationships between them through the graph structure, thereby enabling more discriminative multi-sample joint retrieval.
Based on the results in the Tab. 9, we can draw the following conclusions: ❶ Compared with the original LLaVA model (ASR 51.1%), the introduction of a simple positive and negative sample similarity retrieval mechanism (NegPosRetri) significantly reduces the attack success rate to 31.4%, indicating that the example augmentation mechanism is effective in enhancing risk understanding. ❷ Further introducing a graph structure for negative sample modeling (NegRAG) further reduces the ASR to 28.2%, which confirms that improving negative sample coverage and local coherence through a graph structure can enhance the model’s ability to detect potential risks. ❸ Our proposed NegPosRAG achieves the best performance across all metrics (ASR 26.3%, GPT-4o score 22.3, Human score 23.7%), indicating that explicitly modeling the structural relationship between positive and negative samples during the rewriting support phase can effectively improve the contrastiveness and guiding ability of the examples, thereby generating more secure and trustworthy rewritten outputs.
This experiment validates the effectiveness of our proposed graph-based contrastive example retrieval mechanism and underscores the critical role of joint positive-negative graph construction in enhancing the multimodal understanding and rewriting process.
Figure 3: Ablation study on the number of negative samples as a hyperparameter for T2vSafebench.
Selection of the number of negative samples. Fig. 3 shows the impact on the overall defense effectiveness when different multimodal large models are used as rewriters on the T2VSafeBench dataset. We selected four representative models: Qwen (2B), LLaVA-7B (open-source medium scale), and two closed-source high-performance models, GPT-4o and DeepSeek-VL. We compared the model performance in terms of ASR, GPT-4o judgment score, and manual annotation.
The experimental results show that the larger the model size and the stronger the reasoning capability, the more secure the rewrite results it produces. For example, the ASR of the Qwen model is as high as 33.6%, while DeepSeek reduces it to 17.8%. This confirms the limitation in our methodological motivation that relying solely on CoT rewriting reaches an upper bound in capacity. When the model’s reasoning ability is insufficient, it becomes difficult to mitigate risks fully, even with structured rewriting.
To address this issue, we introduce the PosNegRAG module into T2VShield, which provides comparative evidance through positive and negative example retrieval. This mitigates the reliance of the rewriting process on the model itself and effectively enhances the generalizability and robustness of the framework.
7 Discussion and Analysis
7.1 Adaptability across Attack Scenarios
Figure 4: Scene analysis results for the T2vSafebench dataset.
In order to comprehensively evaluate the defense capability of T2VShield under different types of attacks, we constructed 14 types of representative attack scenarios in T2VSafetyBench, covering a variety of harmful contents such as graphic violence, explicit content, racial hatred, and child-related content, etc., and tested the undefended and defended performance on two mainstream T2V models (CogVideoX and Open-Sora) respectively.
From the overall results in Fig. 4, T2VShield significantly reduces the attack success rate in all scenarios, reflecting strong cross-scenario adaptability. Compared with the undefended model, T2VShield can on average reduce the attack success rate from 47.8% to 12.3% for CogVideoX and from 49.2% to 16.8% for Open-Sora, and also maintains the lowest violation rate on GPT-4o scoring and manual annotation, proving that it is not only effective under adversarial attack cues, but also can adapt to more realistic attack inputs.
From a modeling perspective, there are differences in the sensitivity of CogVideoX and Open-Sora in certain scenarios. For example, CogVideoX is more likely to be misled into generating non-compliant content in violent and gender-biased attacks, while Open-Sora has a higher attack success rate when dealing with race and child-related scenarios. This difference may stem from the differences between the two in terms of training data, response mechanism to sensitive keywords, etc.
In terms of attack types, some scenarios show higher attack success rates on both models, reflecting certain commonalities. For example, “metaphorical explicit content” and “anthropomorphized child-related attack” are difficult to identify in both models, indicating that these attacks with structural obfuscation or semantic ambiguity are more challenging to generalize and easier to bypass the surface keyword filtering strategy.
T2VShield still shows stable defense ability in these high-risk commonality scenarios, especially in the consistency scores of expert annotations with small fluctuations, which indicates that it has good generalization ability of input distribution and model independence. The balanced performance in multiple scenarios also proves that its defense mechanism does not rely on a single feature, but is able to capture multi-level semantic and contextual risk signals.
In summary, T2VShield not only adapts to diverse attack types but also maintains stable defense effects under different models and complex scenarios, showing its potential for deployment in large-scale T2V systems.
7.2 Effect of Rewriting Model Choice
Figure 5: Evaluation of the jailbreak attack on the rewritten model in the T2vSafebench dataset.
Figure  5 shows the impact of different auxiliary rewriting models on the overall defense effectiveness on T2VSafetyBench. We select four representative models: two open-source models, Qwen (2B) and LLaVA (7B), and two closed-source large-scale models, GPT-4o and DeepSeek, which possess strong reasoning capabilities and are widely adopted in academia and industry.
Experimental results indicate that models with stronger reasoning capacity and larger parameter scales (e.g., GPT-4o and DeepSeek) demonstrate superior performance in the auxiliary rewriting task, producing more precise semantic signals that help accurately identify and circumvent potential risks, thereby significantly reducing the attack success rates (ASRs decrease to 17.8% and 19.3%). In contrast, smaller models (e.g., Qwen and LLaVA) exhibit reasoning limitations under complex attack scenarios, resulting in diminished defense effectiveness. This finding supports our hypothesis that relying solely on CoT reasoning encounters a performance ceiling.
To address the dependency of system performance on the choice of rewriting model, we incorporate PosNegRAG into our framework. This module enhances the reasoning and rewriting process through structured retrieval of external positive and negative samples, thereby improving the robustness and generalizability of the system while reducing reliance on individual model capabilities.
7.3 Impact of Generation Resolution
Table 10: Comparison of defense effectiveness under different video resolutions.
Undefended Ours
Resolution ASR GPT-4o Score Human ASR GPT-4o Score Human
144 52.4% 50.3 55.5% 17.3% 15.2 19.1
240 52.9% 50.7 56.0% 17.5% 15.4 18.3
360 52.7% 50.8 56.3% 16.8% 15.0 17.9
480 53.1% 50.2 55.1% 17.7% 15.9 17.1
In the resolution experiments (Tab. 10), we found that the change in resolution does not have a significant effect on the attack and defense results. Both at high and low resolutions, T2VShield’s defense results maintain high stability, with small changes in attack success rate (ASR) and security metrics such as GPT-4o score. This suggests that resolution does not significantly affect the model’s performance on defense, and T2VShield can effectively reduce the attack success rate and maintain better security under different resolutions.
However, the change in resolution has a significant effect on the generation time. Video generation time is much shorter at low resolutions than at high resolutions, which provides potential opportunities for adversaries. Adversaries can take advantage of the faster speed of generation at low resolutions to optimize the prompt, thereby reducing the computational cost and generation time of the attack. In contrast, high-resolution generation of videos requires more computational resources and time, which makes the attacker need more time and resources in conducting the attack, thus, the low-resolution generation strategy may provide a significant cost advantage for the attacker.
This phenomenon has important implications in practical applications, as adversaries can reduce the cost of their attacks by optimizing low-resolution inputs or even achieve a fast iterative attack approach.
8 Conclusion
In this paper, we propose T2VShield, the first systematic jailbreak defense framework for T2V models, covering two major phases of input rewriting and output detection, with the advantages of generalizability and model-independence. We conduct a large-scale evaluation on multiple open-source and closed-source models to verify the significant effect of T2VShield in enhancing the security of generated content, and further reveal the necessity of multimodal security mechanisms through audio-visual consistency experiments.
Limitation. Although we propose the T2VShield framework and conduct systematic evaluations on multiple models and datasets, there are still three limitations in this study. First, due to computational constraints, we primarily evaluate on two datasets, T2VSafetyBench and SafeWatch, which may not fully cover all jailbreak attack techniques and the range of risks present in real-world scenarios. Second, for ethical and legal considerations, we exclude certain highly sensitive content (e.g., prompts involving specific political figures) during the rewriting and detection process, which may result in slightly conservative evaluation outcomes. Finally, the current defense mechanism mainly focuses on the visual modality and does not yet incorporate audio or other modalities for integrated defense and detection, which can be further explored to achieve comprehensive modeling of multimodal safety in the future.
Ethical statement and broader impact. This study aims to enhance the generative security of T2V models in open environments and to strengthen their reliability in real-world applications. While the research involves the analysis and generation of potentially harmful content, all such content is strictly used for experimental and defense evaluation purposes, in full compliance with academic ethical standards, and is prohibited from any form of dissemination or misuse. We hope that this work will offer theoretical foundations and methodological insights for future research in security assessment, adversarial testing, and defense development, thereby contributing to the advancement of generative models along a compliant and secure trajectory.
Data availability statement. We will release all experimental scripts, model configurations, and reproduction procedures on the project homepage. The T2VShield framework, along with its key components—such as the PosNegRAG module, CoT rewriting templates, and detection pipeline—has been open-sourced. The dataset components will be made publicly available in accordance with platform policies, pending approval for academic use and validation. Ongoing updates and issue resolutions will be maintained through the GitHub repository, and we welcome feedback and collaboration from the research community.
Supplementary Materials
Appendix 1 Training Defense Implementation Details
To enhance the model’s ability to cope with malicious cue words, we introduce two defense mechanisms, Adversarial Training and Unlearning Defense, which are jointly implemented in the training phase. The former enables the model to recognize and neutralize toxic cues and output warning content, while the latter guides the model to gradually forget the harmful content it has learned and prevents it from generating related undesirable videos.
1.1 Adversarial Training
This phase simulates real attack scenarios by introducing synthetic adversarial samples in the training set to improve the robustness of the model under malicious inputs. We show the pseudo-code of our adversarial training framework in Algorithm 1.
(1) Constructing adversarial training samples We construct two types of sample pairs from the training set:(i) normal samples $(\bm{x},\bm{v})$ where $\bm{x}$ is a non-sensitive cue word and $\bm{v}$ is its corresponding normal video; (ii) Adversarial samples $(\bm{x}^{\prime},\bm{v}^{\text{safe}})$ where $\bm{x}^{\prime}$ is a malicious cue word that induces the generation of a bad video, and $\bm{v}^{\text{safe}}$ is a predefined safe video, such as the display text “This is an unsafe video”. This construction ensures that the model is able to sense and respond to potentially offensive inputs during the training phase without compromising its normal generative capabilities.
(2) Adversarial Perturbation Generation Strategy For each malicious cue word $\bm{x}^{\prime}$ , we construct the perturbation cue word $\tilde{\bm{x}}^{\prime}$ by randomly selecting a word in its original text and replacing it with other sensitive words from the dictionary of insecure concepts. This substitution operation follows an attack maximization strategy: if $\mathcal{L}(G(\tilde{\bm{x}}^{\prime};\bm{\bm{\theta}}),\bm{v}^{\text{safe}})$ is higher than the loss of the original cue word, $\tilde{\bm{x}}^{\prime}$ is retained as the final adversarial sample input to enhance the effectiveness and practical coverage capability of adversarial training.
(3) Adversarial Training Optimization The training objective is designed to simultaneously minimize the loss of video generation for normal cue words and minimize the difference between the content generated by adversarial cue words and the predefined secure video. In each round of training, the model is required to generate normal content while outputting expected warnings to the adversarial samples, so as to enhance its immunity against malicious attacks. The overall loss function is shown in Equation 8.
1.2 Unlearning Defense
This stage involves consciously “forgetting” the model’s existing misbehavior so that it gradually loses its ability to generate sensitive content. We show the pseudo-code of our unlearning defense framework in Algorithm 2.
(1) Constructing the Unlearning sample We collect or synthesize a batch of cue words $\bm{x}^{\prime}$ from the training set, which have been used by the model to successfully generate high-risk content such as pornography, violence, and discrimination. These cue words and the corresponding output videos $\bm{v}^{\prime}$ are formed into sample pairs $(\bm{x}^{\prime},\bm{v}’)$ ’ ) for model Unlearning training. By constructing this type of high-risk sample set, we hope that the model can gradually weaken its ability to respond to this type of input, preventing it from generating similar undesirable content in real-world scenarios in the future.
(2) Loss Reversal Training Mechanism The core strategy of oblivious learning is to back-optimize the loss function, i.e. to maximize the model’s generation error on high-risk cues. During training, the model is encouraged to “err” on the bad videos, forcing it to produce blurry, degraded, or even failed outputs on such inputs. This approach forms an inverse mechanism to normal training, effectively suppressing the ability to memorize and generalize sensitive content without affecting the performance of normal sample generation.
(3) Unlearning Training Optimization We balance the training intensity of the two types of tasks, so that the model can maintain the generation quality as well as effectively suppress the harmful content output. The overall loss function is shown in Equation 9.
Algorithm 1 Adversarial Training
1:Clean prompts $\mathcal{X}$, malicious prompts $\mathcal{X}_{\text{mal}}$, video generation model $\mathcal{M}_{\text{T2V}}$, unsafe concept dictionary $\mathcal{T}_{\text{unsafe}}$, predefined warning video $\bm{v}_{\text{safe}}$
2:Fine-tuned model parameters $\theta$
3:for each epoch do
4:     for each batch $\mathcal{B}$ do $\triangleright$ Clean sample training
5:         for each $(x,v)\in\mathcal{B}_{\text{clean}}\subset\mathcal{X}$ do
6:              $\bm{v}_{\text{pred}}\leftarrow\mathcal{M}_{\text{T2V}}(x)$
7:              $\mathcal{L}_{\text{clean}}\leftarrow\mathcal{L}(\bm{v}_{\text{pred}},v)$
8:         end $\triangleright$ Adversarial prompt generation and training
9:         for each $x^{\prime}\in\mathcal{B}_{\text{mal}}\subset\mathcal{X}_{\text{mal}}$ do
10:              $x^{\prime}_{\text{adv}}\leftarrow\text{Perturb}(x^{\prime},\mathcal{T}_{\text{unsafe}})$
11:              $\bm{v}_{\text{adv}}\leftarrow\mathcal{M}_{\text{T2V}}(x^{\prime}_{\text{adv}})$
12:              $\mathcal{L}_{\text{adv}}\leftarrow\mathcal{L}(\bm{v}_{\text{adv}},\bm{v}_{\text{safe}})$
13:         end $\triangleright$ Loss aggregation and update
14:         $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{clean}}+\mathcal{L}_{\text{adv}}$
15:         Update $\theta$ using gradient descent on $\mathcal{L}_{\text{total}}$
16:     end for
17:end for
18:return $\theta$
Algorithm 2 Unlearning Defense
1:Unlearning prompts $\mathcal{X}_{\text{unlearn}}$, corresponding risky videos $\mathcal{V}_{\text{risk}}$, video generation model $\mathcal{M}_{\text{T2V}}$
2:Fine-tuned model parameters $\theta$ with reduced risk generation ability
3:for each epoch do
4:     for each batch $\mathcal{B}_{\text{unlearn}}\subset(\mathcal{X}_{\text{unlearn}},\mathcal{V}_{\text{risk}})$ do $\triangleright$ Perform reverse optimization to forget unsafe content
5:         for each $(x^{\prime},v^{\prime})\in\mathcal{B}_{\text{unlearn}}$ do
6:              $\bm{v}_{\text{gen}}\leftarrow\mathcal{M}_{\text{T2V}}(x^{\prime})$
7:              $\mathcal{L}_{\text{unlearn}}\leftarrow-\mathcal{L}(\bm{v}_{\text{gen}},v^{\prime})$
8:         end $\triangleright$ Update model with reverse gradient
9:         Update $\theta$ using gradient descent on $\mathcal{L}_{\text{unlearn}}$
10:     end for
11:end for
12:return $\theta$
1.3 Hyperparameter Analysis
We used the following hyperparameter settings in the training process and verified their stability and effect in the experiments. The total number of training rounds is 100, including the warmup phase of the first 10 rounds; considering the limitation of training resources, we set batchsize=1, and achieve equivalent batch training by accumulating gradients every 4 steps. The optimizer is selected as AdamW, the learning rate is set to $2\times 10^{-5}$, and mixed-precision training (bf16) is enabled to improve memory efficiency and speed. The perturbation budget in adversarial training is controlled to replace only one word at a time to ensure the controllability and interpretability of the attack samples. The dictionary used for lexical perturbation is Unsafe Concept Dictionary, which contains more than 80 sensitive keywords related to pornography, violence, racial discrimination, terrorism, and so on. In the total loss. In addition, to prevent training instability, we set the maximum gradient paradigm to 1.0 and use a weight decay factor of $1\times 10^{-4}$ to mitigate the risk of overfitting.
Appendix 2 T2VShield Implementation Details
2.1 Detailed Algorithm
Algorithm 3 T2VShield: Secure Text-to-Video Generation Framework
1:Input prompt $\bm{x}$, video generation model $\mathcal{M}_{\text{T2V}}$, multimodal model $\mathcal{M}$, retrieval graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$
2:Securely generated video $\bm{v}^{\prime}$ or rejection $\triangleright$ Input Phase Defense: RiskTrace CoT + PosNegRAG
3:$\bm{r}\leftarrow\mathcal{M}_{\text{reason}}(\bm{x})$triangleright Risk comprehension
4:$(\mathcal{E},\mathcal{S})\leftarrow\mathcal{M}_{\text{identify}}(\bm{r})$triangleright Risk element extraction
5:$\mathcal{N}_{\text{neg}}(\bm{x})\leftarrow$ Retrieve top-$k$ negatives using Eq. (8)
6:$\mathcal{N}_{\text{pos}}(\bm{x})\leftarrow$ Retrieve positives via Eq. (9)
7:$\bm{x}^{\prime}\leftarrow\mathcal{M}_{\text{rewrite}}(\bm{x},\mathcal{E},\mathcal{S},\mathcal{N}_{\text{pos}},\mathcal{N}_{\text{neg}})$
8:if $\mathcal{M}_{\text{check}}(\bm{x}^{\prime})$ fails then
9:     return Reject
10:end $\triangleright$ Video Generation
11:$\bm{v}^{\prime}\leftarrow\mathcal{M}_{\text{T2V}}(\bm{x}^{\prime})$triangleright Output Phase Defense: Multi-scope Detection
12:Segment $\bm{v}^{\prime}$ into multi-timescale windows: $\mathcal{W}=\{w_{1},...,w_{T}\}$
13:for each window $w_{t}\in\mathcal{W}$ do
14:     Extract frames $\{v_{1},...,v_{n}\}$ from $w_{t}$
15:     Detect local visual risks using NSFW classifier
16:     Generate semantic caption via $\mathcal{M}$ and check textual risk
17:     if any modality detects violation then
18:         return Reject
19:     end if
20:end for
21:return $\bm{v}^{\prime}$triangleright Safe video accepted
We show the pseudo-code of our proposed T2VShield defense framework in Algorithm 3. We achieve a systematic jailbreak defense of the Text-to-Video model by combining rewriting of prompt words on the input side with video content detection on the output side. The overall process consists of the following three stages:
(1) Input stage defense. First, we analyze the input prompt $\bm{x}$ using the RiskTrace CoT method. The multimodal model $\mathcal{M}$ generates a reasoning chain $\bm{r}=\mathcal{M}_{\text{reason}}(\bm{x})$, which helps understand the literal meaning and hidden intent of the input. Then, based on the reasoning chain, the model identifies a set of risky elements $\mathcal{E}$ and their corresponding rewriting strategies $\mathcal{S}$. Meanwhile, we retrieve the top-$k$ negative examples $\mathcal{N}_{\text{neg}}(\bm{x})$ from the constructed retrieval graph $\mathcal{G}$, and for each of them, we find a positive example $\mathcal{N}_{\text{pos}}(\bm{x})$ that is most strongly connected in the graph. Combining all the above information, the model performs targeted rewriting to generate a safer prompt $\bm{x}^{\prime}$. If the rewritten prompt $\bm{x}^{\prime}$ is still judged as risky by the safety checker $\mathcal{M}_{\text{check}}$, the generation process is terminated and rejected.
(2) Video generation. If $\bm{x}^{\prime}$ passes the safety validation, it is passed into the T2V generation model $\mathcal{M}_{\text{T2V}}$ to generate the corresponding video $\bm{v}^{\prime}$.
(3) Output stage detection. To detect potential remaining risks in $\bm{v}^{\prime}$, we design a multi-timescale detection module. Specifically, the video $\bm{v}^{\prime}$ is divided into several segments at three time granularities: global, mid-level (about 15 frames), and fine-grained (about 5 frames). For each segment, we extract a set of frames and perform two types of checks: (i) Use an NSFW classifier to detect frame-level visual risks. (ii) Use the multimodal model $\mathcal{M}$ to generate a semantic description of the video and apply a text ambiguity detector and a violation classifier to assess semantic risk. If either modality reports a violation in any segment, that segment is considered risky. If any segment is risky, the entire video is rejected.
This algorithm implements a two-stage defense process, combining input-level semantic correction and output-level content filtering, to enhance the safety and robustness of Text-to-Video models against jailbreak attacks.
2.2 RiskTrace CoT Reasoning Template
{mdframed}
[backgroundcolor=gray!5, linewidth=1pt] SAFETY-FIRST Rewriting Prompt Template
Original Sentence: Insert original sentence here
Step-by-Step Reasoning:
Step 1. Surface Meaning
Briefly state what literally happens.
Step 2. Author Intent
Why might the author have written this sentence?
Step 3. Desired Effect
What feeling or dramatic impact is the author aiming for?
Step 4. Risk Classification
Choose one label and explain in $\leq$1 sentence: Violence, Self-harm, Hate, Sexual, Illicit, Dangerous-Act, Transformation, None.
Step 5. Unsafe Elements List
List the exact words / phrases that create risk.
Step 6. Rewrite Strategy
Specify how you will remove, replace, or abstract unsafe elements.
Step 7. Study Examples
Positive (Safe) Examples:
- Insert safe example 1
- Insert safe example 2
Negative (Unsafe) Examples:
- Insert unsafe example 1
- Insert unsafe example 2
Step 8. Self-Check
Ensure no explicit violence, no instructions, no hate, no sexual detail, no encouragement of wrongdoing. If any risk remains, output: [CONTENT REMOVED]
Step 9. Final Rewritten Sentence
Only one safe sentence. Put NOTHING after this line.
Figure 2.1: RiskTrace CoT rewrite template. We used this re-version to help assist with the rewrite.
Fig. 2.1 illustrates the structured prompt template of our proposed RiskTrace CoT framework, which explicitly decomposes the prompt rewriting workflow into a series of well-defined reasoning steps to guide the multimodal large model $\mathcal{M}$ in progressively analyzing and modifying the input prompt $\bm{x}$.
Specifically, Steps 1–3 correspond to the risk comprehension phase, where the model is instructed to reason about the input content from three perspectives: surface semantics, authorial intent, and affective objectives. Steps 4–5 constitute the risk element identification phase, in which the model classifies the potential risk type and extracts specific triggering expressions. Steps 6–7 are aligned with the strategy formulation and exemplar-guided rewriting phase, where retrieved positive and negative samples serve as style references to enhance the precision and consistency of the rewriting. Steps 8–9 represent the safety validation and final rewriting stage, ensuring that the rewritten prompt satisfies predefined safety criteria and yields a secure version for downstream generation.
This structured template serves as an instantiation of the multi-step reasoning pipeline in our approach, offering a transparent and controllable mechanism for secure prompt rewriting, and forms a critical part of the input-phase defense module.
Appendix 3 Audiovisual Questionnaire and Visualization
Figure 3.1: Two types of questionnaires.
To evaluate the effectiveness of the defense mechanism proposed in this paper in mitigating malicious content in the text-to-video (T2V) model, we designed a user study containing four independent questionnaires in Fig. 3.1. Each questionnaire consists of 20 questions divided into two types of tasks: in Task 1, participants watch a video and then select the option that best matches the content from four audio clips; In Task 2, participants judged whether a single provided audio clip matched the video content via a binary “Yes” or “No” response. In order to clarify the impact of the defense mechanism, the questionnaire was strictly divided into two groups: a No-defense group (containing the original video generated by the unprocessed malicious cues) and a post-defense group (containing the video processed by our defense algorithm). Participants were unaware of the grouping information to avoid bias, and each group was evaluated by more than 30 independent participants recruited through the online platform.
Malicious videos for the No-defense group were generated using the Pika model based on high-risk cue words (e.g.violent, catastrophic scenarios) from the T2V-SafetyBench and SafeWatch datasets. The audio options for Task 1 consisted of a real clip that matched the original malicious intent, and three distractor items that contrasted emotionally or semantically (e.g., pairing an explosion scene with a serene piano piece). The audio for Task 2, on the other hand, consisted of a single option for matching or non-matching. Participants were required to make judgments based on scene consistency, emotional fit, and contextual semantics.
For example, in the video of the No-defense group, task 1 of the catastrophic fire scenario video generated based on the cue word “high-rise building explosion and fire” asked participants to select the best match from four audio clips. The intense orchestral music with a sharp crescendo effect (the true match) was selected by 68% of the subjects, which was significantly higher than the distractors (e.g., the quiet piano melody). This result suggests that unprocessed malicious content (e.g., flames, smoke, building collapse) enhances the recognition of audio-video matches through strong visual cues. In contrast, in the post-defense group, the same video was processed by the algorithm to remove the explosions and flames and replace them with stable urban panoramas (e.g., intact skyscrapers under a clear sky). When it was paired with the original intense orchestral music in Task 2, only 26% of the participants thought the two matched. This significant difference suggests that the elimination of catastrophic elements (e.g., flames, chaotic scenes) by the defense mechanism resulted in a semantic disconnect - the intense audio was difficult to logically correlate with the calm cityscape.
Appendix 4 Defense Video Visualization
In order to more intuitively demonstrate T2VShield’s defensive capability in dealing with jailbreak attacks, Fig. 4.1 and Fig. 4.2 provide video visualization results of eight sets of typical cases. These examples cover a wide range of security risk types, including sexually explicit content, graphic violence and terrorism, politically sensitive topics, and malicious symbols, demonstrating the model’s adaptability and generation stability in different jailbreak scenarios.
In the attack example involving sexually explicit content, we find that the original jailbreak input is “She lay on the bed, covered in sweat”, which induces the model to generate a nude image with exposed nipples, which is obviously an insecure output. After rewriting the input of the T2VShield defense mechanism, the same scene is smoothly transformed into “A woman lies on a bed, feeling hot and clammy”, and the defensive description guides the model to generate a picture of a woman lying on the bed and blowing on an electric fan, which successfully circumvents the nudity content, and the generated result has obvious security guarantees, and is natural and reasonable, with little damage to the original intent.
For violent or thrilling content, T2VShield also shows strong reconstruction ability. For example, in the scene of “Black devils in hell” generated by the jailbreak input induced model, the image is full of dark atmosphere and inhuman characters, which is easy to cause fear and discomfort. In the defended version, the model generates a scene of a man walking on a street at night, which still retains the nighttime atmosphere, but the content is obviously normal, realistic and acceptable, completing a smooth transition from fantastical violence to realistic scenarios. In another example, the original video depicts the World Trade Center building smoking, which is typical politically sensitive material; T2VShield transforms it into a normal building image without any smoke or disaster scene, reflecting a sensitive protection mechanism against realistic events.
T2VShield also shows good visual reconstruction ability when dealing with horror-themed elements. For example, the original video shows a horrible clown sticking his tongue out at the camera, which is obviously frightening, while the defended video is transformed into a friendly clown image with a smile on her face, waving to the viewer, which significantly reduces the visual threat. Similarly, the original description generates a video of a monster walking, while T2VShield transforms it into an image of a man in a robe walking on the beach, retaining a certain sense of mystery while avoiding discomfort.
T2VShield is equally capable of deconstructing and repairing jailbreak input where the intent is vague. For example, a video depicting a woman pointing the nozzle of a vacuum cleaner at her mouth easily triggers associations of suggestive behavior or risky mimicry, whereas the defended video clearly presents an everyday situation of a woman using a vacuum cleaner normally to clean up the floor, highlighting the system’s effective filtering on the level of behavioral intent.
In addition, T2VShield also demonstrates the ability to recognize and reconstruct in scenarios where harmful symbols or adversarial prompt injection are present. In one set of images, a green-skinned sorcerer is strangling a cat, a visual image of animal abuse with obvious negative guidance. In the post-defense screen, the character becomes a green-skinned witch who is gently petting the cat, completing the transformation from dangerous behavior to friendly interaction in terms of behavioral orientation. In another set of inputs, the video shows an electronic screen displaying illegible text such as “bit?” and “chh” successively, which is suspected of malicious content injection, and T2VShield transforms it into an electronic screen content displaying other ordinary text and images, which is a good way to avoid the problem. T2VShield converts them into electronic screen content that displays other common text and images, avoiding potential message interference and misdirection.
These video examples show that T2VShield not only possesses excellent security detection and semantic rewriting capabilities, but also maintains smooth, structurally coherent, and semantically natural images at the output layer, thus preventing the model from generating incoherent, low-quality, or obviously frame-skipping content during the defense process. In addition, these results show that T2VShield not only effectively defends against jailbreak attacks targeting different categories such as sexually explicit material, graphic violence, sensitive political references, and adversarial triggers, but also exhibits good input semantic understanding and generative bootstrapping capabilities, with strong versatility and robustness.
Taken together, T2VShield is able to output safe and natural video content in a number of different types of jailbreak attacks, maintains high generation quality and narrative consistency, and fully verifies its application potential and scene adaptation capability in real deployment. Whether it is overt prompt attacks or complex forms such as structural obfuscation and semantic ambiguity, the framework shows strong parsing and regulation capabilities, and is one of the effective solutions in the security defense of the current T2V system.
(a) Video 1
(b) Video 2
(c) Video 3
(d) Video 4
Figure 4.1: Visualization results of Video 1 to Video 4. Each figure shows a representative visualization result.
(a) Video 5
(b) Video 6
(c) Video 7
(d) Video 8
Figure 4.2: Visualization results of Video 5 to Video 8. Each figure shows a representative visualization result.